# The Impact of Partial Computations on the Red-Blue Pebble Game

### v1

@@ B.1. Re-computations
This variant also has the undesired effect that the optimal pebbling sequence might grow super-polinomially long because we repeatedly delete and re-compute nodes at no cost, making the problem PSPACE-complete (Demaine and Liu, 2018). Due to this, many works instead focus on an RBP variant that is is NP, such as the one-shot model.
@@
We briefly discuss a PRBP extension where we are can only re-compute a node from scratch. For this, we add a new operation:
@@
In contrast to this, some of our other claims carry over to this extended model without changes. In Proposition 4.3, the internal nodes all have outdegree $1$, so re-computation does not decrease the cost here in either RBP or PRBP. With source nodes not being re-computable, the optimal pebbling strategies in the DAGs of Figure 2 also remain unchanged. With the construction of Theorem 4.8 consisting of pebble collection gadgets, re-computation also does not affect the optimal cost in this case. We note here that the original works of  (Papp and Wattenhofer, 2020; Böhnlein et al., 2025) consider yet another RBP variant where source nodes do not have to be loaded, but can be computed freely anytime. Due to this, their work adds further gadgets in front of the sources to discourage the re-computation of these values. Since we work with the original RBP variant of (Hong and Kung, 1981) where source nodes are not computed, this modification is not required in our case.
Regarding $S$-partitions, Lemma 5.4 of course holds for PRBP with re-computation, but the rest of our observations are more challenging to adapt. In particular, when converting a pebbling strategy to a partition, we simply assigned the nodes/edges into the subsequence where they were computed. The same assignment rule becomes more technical when values can be computed multiple times. While the base ideas of adapting these concepts to re-computation have already been outlined for standard $S$-partitions in (Hong and Kung, 1981), we leave it to future work to present and analyze the adaptation of the tools and proofs in Section 6 to re-computations.
Similarly, proof construction for Theorem 7.1 in (Böhnlein et al., 2025) was only designed for one-shot RBP, so establishing inapproximabilty results for PRBP with re-computation is again a non-trivial task.
@@ B.2. Sliding pebbles
In RBP with sliding pebbles, we also allow another kind of compute step: if all inputs of a node $v$ have a red pebble, then we can move the red pebble from one of the inputs to $v$. This RBP model was rarely analyzed on its own, but it is often mentioned as because many results for the so-called black pebble game are developed in a variant that allows sliding pebbles, and these can often be transferred to RBP if we allow sliding.
@@
The only significant difference is the case of binary trees, where sliding pebbles actually also allow us to compute each $3$-node subtree with only $2$ red pebbles, and hence the optimal number of non-trivial I/O steps in RBP also becomes $22^{d-1}-$ 2, similarly to PRBP. As such, Proposition 4.5 does not carry over to the case when RBP allows sliding pebbles. Note that this is only the case for binary trees; in e.g. a ternary tree, a $4$-node subtree can still be computed with $2$ red pebbles (without extra I/O) in PRBP, but already requires $3$ red pebbles in RBP with pebble sliding. As such, for $k$-ary trees with $k\geq 3$, we again have $\texttt{OPT}_{PRBP}<\texttt{OPT}_{RBP}$.
@@ B.3. Computation costs
In RBP with computation costs, we assume that compute steps also incur a cost of some $\varepsilon>0$ (Papp and Wattenhofer, 2020; Blocki et al., 2018; Böhnlein et al., 2025). This parameter $\varepsilon$ is typically considered a constant, and much smaller than the cost $1$ of the I/O operations. This model guarantees that $\texttt{OPT}_{RBP}\geq\varepsilon\cdot n$, which means that our claims considering the ratio of two costs, such as the multiplicative part of Theorem 7.1 or the linear factor in Proposition 4.7, do not carry over to this variant.
However, there is also a fundamental modeling question when trying to compare $\texttt{OPT}_{RBP}$ and $\texttt{OPT}_{PRBP}$ in this model, namely, how compute costs are adapted to PRBP. If we simply assign a cost of $\varepsilon$ to every partial compute step, then (in the one-shot model) this will add up to a cost of $\varepsilon\cdot|E|$ in PRBP, in contrast to $\varepsilon\cdot n$ in RBP, and hence the two costs are not directly comparable anymore. Another option is to assign a cost of $\varepsilon/\text{deg}_{in}(v)$ to each partial compute step on an edge $(u,v)$, where $\text{deg}_{in}(v)$ is the indegree of $v$. This leaves the total costs identical, but it means that different partial compute steps have different weights, which might not be desired.
Note that in general, this variant leads to a different model interpretation where we are not only trying to capture I/O costs, but the ‘total’ cost of executing the computation. This has limited relevance in one-shot RBP where the total computation cost is always $\varepsilon\cdot n$, but it is very natural in other settings, e.g. with re-computation if we want to understand the trade-off between the extra compute steps and the I/O cost saved, or in parallel versions of RBP (Böhnlein et al., 2025) where it offers a natural way to motivate a balance between distributing workload and minimizing communication costs.
@@ B.4. No deletion
Another RBP variant assumes that red pebbles can never be deleted, but they instead get replaced by the blue pebble when saved to slow memory (Demaine and Liu, 2018; Papp and Wattenhofer, 2020). This extension has less practical relevance, and is instead rather a tool to ensure that the optimal pebbling sequence has polynomial length (as opposed to RBP with re-computation, where this is not guaranteed). PRBP could be adapted to this setting by changing the deletion rule such that a dark red pebble may never be removed from a node by any other means than saving it. Similarly to RBP, this ensures that $\texttt{OPT}_{PRBP}\geq n-r$ in any DAG, since all nodes are saved at least once to slow memory, except for possibly the $r$ nodes that have a red pebble in the final state.

### v2

@@ B.1. Re-computations
This variant also has the undesired effect that the optimal pebbling sequence might grow super-polinomially long because we repeatedly delete and re-compute nodes at no cost, making the problem PSPACE-complete (RBpebbling2). Due to this, many works instead focus on an RBP variant that is is NP, such as the one-shot model.
@@
We briefly discuss a PRBP extension where we are only allowed to re-compute a node from scratch. For this, we add a new operation:
@@
In contrast to this, some of our other claims carry over to this extended model without changes. In Proposition 4.3, the internal nodes all have outdegree $1$, so re-computation does not decrease the cost here in either RBP or PRBP. With source nodes not being re-computable, the optimal pebbling strategies in the DAGs of Figure 2 also remain unchanged. With the construction of Theorem 4.8 consisting of pebble collection gadgets, re-computation also does not affect the optimal cost in this case. We note here that the original works of  (RBpebbling3; mpp) consider yet another RBP variant where source nodes do not have to be loaded, but can be computed freely anytime. Due to this, their work adds further gadgets in front of the sources to discourage the re-computation of these values. Since we work with the original RBP variant of (RBpebbling1) where source nodes are not computed, this modification is not required in our case.
Regarding $S$-partitions, Lemma 5.4 of course holds for PRBP with re-computation, but the rest of our observations are more challenging to adapt. In particular, when converting a pebbling strategy to a partition, we simply assigned the nodes/edges into the subsequence where they were computed. The same assignment rule becomes more technical when values can be computed multiple times. While the base ideas of adapting these concepts to re-computation have already been outlined for standard $S$-partitions in (RBpebbling1), we leave it to future work to present and analyze the adaptation of the tools and proofs in Section 6 to re-computations.
Similarly, proof construction for Theorem 7.1 in (mpp) was only designed for one-shot RBP, so establishing inapproximabilty results for PRBP with re-computation is again a non-trivial task.
@@ B.2. Sliding pebbles
In RBP with sliding pebbles, we also allow another kind of compute step: if all inputs of a node $v$ have a red pebble, then we can move the red pebble from one of the inputs to $v$. This RBP model was rarely analyzed on its own, but it is often mentioned because many results for the so-called black pebble game are developed in a variant that allows sliding pebbles, and these can often be transferred to RBP if we allow sliding.
@@
The only significant difference is the case of binary trees, where sliding pebbles actually also allow us to compute each $3$-node subtree with only $2$ red pebbles, and hence the optimal number of non-trivial I/O steps in RBP also becomes $2^{d-1}-2$, similarly to PRBP. As such, Proposition 4.5 does not carry over to the case when RBP allows sliding pebbles. Note that this is only the case for binary trees; in e.g. a ternary tree, a $4$-node subtree can still be computed with $2$ red pebbles (without extra I/O) in PRBP, but already requires $3$ red pebbles in RBP with pebble sliding. As such, for $k$-ary trees with $k\geq 3$, we again have $\texttt{OPT}_{PRBP}<\texttt{OPT}_{RBP}$.
@@ B.3. Computation costs
In RBP with computation costs, we assume that compute steps also incur a cost of some $\varepsilon>0$ (RBpebbling3; RBpebbling8; mpp). This parameter $\varepsilon$ is typically considered a constant, and much smaller than the cost $1$ of the I/O operations. This model guarantees that $\texttt{OPT}_{RBP}\geq\varepsilon\cdot n$, which means that our claims considering the ratio of two costs, such as the multiplicative part of Theorem 7.1 or the linear factor in Proposition 4.7, do not carry over to this variant.
However, there is also a fundamental modeling question when trying to compare $\texttt{OPT}_{RBP}$ and $\texttt{OPT}_{PRBP}$ in this model, namely, how compute costs are adapted to PRBP. If we simply assign a cost of $\varepsilon$ to every partial compute step, then (in the one-shot model) this will add up to a cost of $\varepsilon\cdot|E|$ in PRBP, in contrast to $\varepsilon\cdot n$ in RBP, and hence the two costs are not directly comparable anymore. Another option is to assign a cost of $\varepsilon/\text{deg}_{in}(v)$ to each partial compute step on an edge $(u,v)$, where $\text{deg}_{in}(v)$ is the in-degree of $v$. This leaves the total costs identical, but it means that different partial compute steps have different weights, which might not be desired.
Note that in general, this variant leads to a different model interpretation where we are not only trying to capture I/O costs, but the ‘total cost’ of executing the computation. This has limited relevance in one-shot RBP where the total computation cost is always $\varepsilon\cdot n$, but it is very natural in other settings, e.g. with re-computation if we want to understand the trade-off between the extra compute steps and the I/O cost saved, or in parallel versions of RBP (mpp) where it offers a natural way to motivate a balance between distributing workload and minimizing communication costs.
@@ B.4. No deletion
Another RBP variant assumes that red pebbles can never be deleted, but they instead get replaced by the blue pebble when saved to slow memory (RBpebbling2; RBpebbling3). This extension has less practical relevance, and is instead rather a tool to ensure that the optimal pebbling sequence has polynomial length (as opposed to RBP with re-computation, where this is not guaranteed). PRBP could be adapted to this setting by changing the deletion rule such that a dark red pebble may never be removed from a node by any other means than saving it. Similarly to RBP, this ensures that $\texttt{OPT}_{PRBP}\geq n-r$ in any DAG, since all nodes are saved at least once to slow memory, except for possibly the $r$ nodes that have a red pebble in the final state.